OpenStudy (anonymous):

Please HELP... MEDAL AND FAN Estimate the value of Square root 52 to the nearest whole number. 9 8 7 6

OpenStudy (anonymous):

@rishavraj

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

Can you guys help me?..

OpenStudy (anonymous):

My answer is 8 I guess

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Or 9.. I'm not sure...

OpenStudy (anonymous):

it will be 7

OpenStudy (anonymous):

How do we do this..

OpenStudy (anonymous):

7*7 = 49- 52 = |-3| 8*8 = 64 - 52 = |12| 3<12

OpenStudy (anonymous):

Oh ok

OpenStudy (uchezzy):

its 7

OpenStudy (anonymous):

hanks

OpenStudy (anonymous):

*Thanks* LOL

OpenStudy (anonymous):

welcome